# In-N-Out: A Parameter-Level API Graph Dataset for Tool Agents

### v1

@@ 4 Experiments
Our experiments address two main questions: (1) can LLMs construct parameter-level API graphs comparable to human annotations, and (2) do such graphs improve downstream tool-agent performance? To answer these, we first benchmark LLMs on graph construction with and without fine-tuning on In-N-Out (§4.1). We then evaluate whether explicit API graphs help agents retrieve prerequisite APIs that supply inputs to another (§4.2), and select interdependent API subsets to support multi-tool query generation (§4.3). We compare using gold graphs versus graphs generated by the fine-tuned LLM, quantifying the remaining gap due to imperfect graph construction.
@@ 4.2 Tool Retrieval with API Graphs
We compare three retrieval settings. In the no-graph setting, all APIs are ranked purely by embedding similarity to the missing parameter description. The gold graph setting uses ground-truth edges from In-N-Out to guide re-ranking. The automated graph setting applies the same re-ranking strategy using edges predicted by the fine-tuned Qwen2.5-32B model analyzed in §4.1, where edge types were inferred for all annotated parameter pairs in the cross-dataset setting.
@@ 4.3 Structured API Subset Selection for Multi-Tool Query Generation
We use GPT-4o mini as the base model and evaluate three conditions from §4.2. In the no-graph setting, the model selects subsets using only API documentation, with no structural guidance. In the gold graph setting, we additionally provide the ground-truth connections between the given APIs as part of the task instruction. In the automated graph setting, the same type of connection information is provided using edges predicted by Qwen2.5-32B under the cross-dataset setup. Prompt details are provided in Appendix E.
@@ 5 Conclusion
Our study also reveals two main limitations. First, In-N-Out currently covers only two benchmark datasets. Expanding to broader domains would improve generalizability—particularly in heterogeneous or rapidly evolving tool ecosystems. Second, our pipeline requires several filtering steps before annotation. Future work may explore semi-automated approaches, such as LLM-assisted labeling with human oversight, to streamline the process while maintaining quality. These directions would further narrow the gap between automated and gold graphs and support more robust tool agent development.

### v2

@@ 4 Experiments
Our experiments address two main questions: (1) can LLMs construct parameter-level API graphs comparable to human annotations, and (2) do such graphs improve downstream tool-agent performance? To answer these, we first benchmark LLMs on graph construction with and without fine-tuning on In-N-Out (§4.1). We then evaluate whether explicit API graphs help agents retrieve prerequisite APIs that supply inputs to another (§4.2), and select interdependent API subsets to support multi-tool query generation (§4.3). We compare using gold graphs versus graphs generated by the fine-tuned LLM, quantifying the remaining gap due to imperfect graph construction. Finally, we conduct an end-to-end evaluation on $\tau$-bench (§4.4) to examine whether the retrieval-level improvements observed in §4.2 translate into better task completion in realistic, multi-turn tool-use settings.
@@ 4.2 Tool Retrieval with API Graphs
We compare three retrieval settings. In the no-graph setting, all APIs are ranked purely by embedding similarity to the missing parameter description. The gold graph setting uses ground-truth edges from In-N-Out to guide re-ranking. The automated graph setting applies the same re-ranking strategy using edges predicted by a Qwen2.5-32B model that is fine-tuned on the other dataset (cross-dataset setup).
@@ 4.3 Structured API Subset Selection for Multi-Tool Query Generation
We use GPT-4o mini as the base model and evaluate three conditions from §4.2. In the no-graph setting, the model selects subsets using only API documentation, with no structural guidance. In the gold graph setting, we additionally provide the ground-truth connections between the given APIs as part of the task instruction. In the automated graph setting, the same type of connection information is provided using edges predicted by the Qwen2.5-32B model fine-tuned on the other dataset (cross-dataset setup). Prompt details are provided in Appendix E.
@@ Results.
4.4 End-to-End Performance Evaluation
While our earlier experiments (§4.2) demonstrate that API graphs substantially improve the retrieval of prerequisite APIs, this raises the question of whether such improvements translate into better end-to-end agent performance. To examine this, we conduct an additional evaluation using the $\tau$-bench Yao et al. (2025), which evaluates multi-turn tool use in realistic user–agent dialogues.
Setup.
$\tau$-bench consists of two domains, $\tau$-retail and $\tau$-airline, containing 15 and 13 API tools, respectively, and 115 and 50 test instances. Since the API documentation in $\tau$-bench does not include output parameter information, we manually created extended documentation by inspecting each API’s invocation results. From these, we enumerated all possible input–output parameter pairs—3,673 for $\tau$-retail and 7,033 for $\tau$-airline.
To construct API graphs, we used the Qwen2.5-32B model fine-tuned on the AppWorld portion of In-N-Out to predict edge types for all parameter pairs and retained only those classified as strong edges. The resulting graphs contained 182 edges for $\tau$-retail and 537 edges for $\tau$-airline. Using these graphs, we augmented each API’s documentation by appending prerequisite API information to the corresponding input parameter descriptions. This allows the agent to reference dependency-aware documentation during tool planning and execution.
Comparison Methods.
Following the experimental setup of $\tau$-bench, we adopted the tool calling agent described in the original paper as our base implementation. This agent uses function calling mode, where the model autonomously decides at each turn whether to respond to the user or invoke a tool. We compare two settings: the baseline, which uses the original $\tau$-bench documentation, and the automated graph setting, where prerequisite API information predicted from our constructed graphs are incorporated into each tool’s documentation. Both GPT-4o-mini and GPT-4o are used as the underlying reasoning models.
Results.
Table 6 summarizes the end-to-end task accuracy across both domains. Overall, incorporating graph information led to consistent improvements over the baseline, with the largest gain observed on $\tau$-airline using GPT-4o (from 50% to 60%). This indicates that structured dependency cues can help higher-capacity models utilize prerequisite information more effectively during tool planning.
Notably, when we qualitatively inspected the agent’s tool-calling trajectories, we observed that the graph-augmented documentation helped the agent correctly identify and invoke prerequisite APIs capable of supplying missing input values for subsequent calls. For instance, in $\tau$-airline baseline setting, the agent originally hallucinated a passenger’s dob (date of birth) by substituting it with the flight’s departure date. After we added to the documentation that get_user_details provides the required dob parameter, the agent correctly called get_user_details first and then used the retrieved date in the booking API, demonstrating the utility of structural guidance.
Model Condition Accuracy (%)
$\tau$-airline $\tau$-retail
GPT-4o-mini Baseline 20.0 37.4
Automated Graph 26.0 40.0
GPT-4o Baseline 50.0 60.9
Automated Graph 60.0 62.6
Table 6: End-to-end task accuracy on $\tau$-bench. Results are reported for the tool calling agent using original documentation (Baseline) versus graph-augmented documentation (Automated Graph).
Discussion.
Although graph augmentation led to measurable improvements, the overall gains were modest—a result that is closely reflects the inherent challenges of $\tau$-bench. According to the failure breakdown reported in the original paper, 56% of errors stem from using wrong arguments or information, 25% from incorrect decisions, and 19% from partially resolved compound requests. Among these, only the first category can be partially addressed by the dependency structure encoded in our API graphs, whereas decision errors or domain-rule violations are largely orthogonal to graph-based reasoning. Moreover, a single task in $\tau$-bench can exhibit multiple failure types simultaneously, making overall end-to-end improvements difficult to achieve.
Nevertheless, these findings suggest that while API graphs alone cannot resolve all reasoning failures, they meaningfully enhance the agent’s ability to provide correct inputs for dependent API calls. Future work may explore more sophisticated ways of integrating API graphs into the agent’s planning and dialogue process, beyond static documentation augmentation.
@@ 5 Conclusion
6 Limitations
Our study has several limitations that suggest directions for future research. First, In-N-Out currently covers only two benchmark datasets. Expanding to broader domains would improve generalizability—particularly in heterogeneous or rapidly evolving tool ecosystems.
Second, some of our filtering and annotation criteria—such as judging whether a connection feels natural—can be inherently subjective. While we aligned annotator criteria through joint calibration, future work could further decompose the notion of naturalness into more objective dimensions. For instance, checking whether a user’s private information is being paired with another person’s public data can help flag unnatural information flow. Likewise, checking whether a parameter is being used across unrelated API domains can reveal domain-inappropriate usage. Such dimensions can provide clearer and more reproducible labeling guidelines.
Finally, our construction pipeline involves multiple filtering stages and substantial manual annotation. As the number of parameter pairs grows, the time and sustained attention required from annotators become a significant bottleneck. Future research could incorporate LLM-assisted labeling—for example, generating example input/output values to help annotators reason more intuitively about parameter compatibility. Complementary strategies such as retrieving similar previously labeled pairs may further support consistent decisions in ambiguous cases. In addition, a continual-learning workflow could reduce manual effort over time: a model trained on In-N-Out could propose labels for new parameter pairs, humans would verify them, and the refined annotations would be incorporated back into training. These directions would enable scalable dataset construction over a larger set of parameter pairs, which in turn may narrow the performance gap between automated and gold graphs and facilitate more robust tool-agent development.